A machine having a V.R 4 needs 2000J of energy to left a load of 400N through a perpendicular distance of 2.5m.Calculate
A.efficiency
B.M.A
1
Expert's answer
2020-10-12T07:47:06-0400

Solution

Given data

Velocity ratio (VR)=4

Load(L)=400N

Distance moved by load=2.5 m

According to velocty ratio

"VR=\frac{displacement\space of\space effort}{displacement\space of\space load}"



So by putting the value


"4=\frac{displacement\space of \space effort}{2.5}"


so

Displacement of effort ="4\times 2.5"

=10 m

And

Work done by Effort=Effort.displacement of Effort

By putting the value

"2000=Effort\cdot10\\Effort=\frac{2000}{10}=200 N"

And

(B)

"MA=\frac{load(L)}{effort(E)}\\MA=\frac{400}{200}=2"


"MA=2"

(A)

"Efficiency=\frac{MA}{VR}"


by putting the value

"Efficiency (\eta)=\frac{2}{4}=\frac{1}{2}"

So

"\eta=50" %


So Mechanical advantage is 2 and efficiency is 50%.